Mary 3 years older than sue, Tom is 5 years older than Mary, the three combine ages are 62. How old is sue
Answer by macston(5194)

Mary 3 years older than sue, Tom is 5 years older than Mary, the three combine ages are 62. How old is sue
S=Sue's age; M=Mary's age=S+3; T=Tom's age=M+5=S+8
S+M+T=62 Substitute for M and T.
S+(S+3)+(S+8)=62
3S+11=62 Subtract 11 from each side.
3S=51 Divide each side by 3.
S=17 ANSWER: Sue is 17 years old.
CHECK:
M=S+3=17+3=20 Mary is 20 years old.
T=M+5=20+5=25 Tom is 25 years old.
S+M+T=62
17+20+25=62
62=62